40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3843     Entry time: Tue Nov 2 00:17:01 2010
Author: Suresh 
Type: Configuration 
Category: Locking 
Subject: Temporary changes to the Video Mux 

Fiber coupling 1064 nm light at the end of X arm

This is 'work in progress'.  The attempt is to bring a few milliwatts of the 1064 nm light from the NPRO at the end of the South(X) Arm to the PSL table through an single mode optical fiber.  This would enable us to tune the two NPRO's to be less than 15 MHz apart by looking at their beat frequency before doubling.  Because we have a 1GHz bandwidth PD at 1064 nm, while the photodiode for green has a BW of about 30MHz.

A PBS (P-type) cube has been introduced into the beam of the X arm NPRO  (between the lamda/2 plate and the input lens of the doubling crystal).  By rotating the face of the PBS slightly away from normal incidence, I have diverted away 1.5mW of the 1064 light for coupling into the fiber. The beam has shifted slightly because of this and the green beam from the south arm has to be realigned to reach the PSL table.

A single mode fiber (Thorlabs SM980-5.8-125),  which was already laid half way, has been extended all the way to the PSL table. It runs along the  South arm in the cable tray. 

A pair of mirrors have been arranged in a zig-zag to steer the beam into a fiber coupler.  There was some hope that this coupler had been aligned at some point in the past and that attaching a fiber might result in some transmission.  But this is not the case and fiber coupler needs to be readjusted.

In order to see the light transmitted through the fiber, a camera has been set up on the PSL table.  Its output has been routed into the 'Ref Cavity reflected' video signal.  A video cable running from the ETMX to the Video-MUX used to be connected to the input channel 9 of the Video MUX.  This has now been shifted to output channel 25 of the MUX and disconnected from the camera at the ETMX.   The 'Ref Cav Refl.' video signal has been routed to the output channel 25.  The camera looking at the fiber output can now be seen on a local monitor at the end of the X arm and on the video monitor in the control room.

With the fiber disconnected, the 1064 nm beam was steered into the fiber coupler and its transmission maximised by observing  with an IR viewer.  The fiber was then connected and then the transmission at the PSL table was sought.  There was no transmission seen after a searching around this region for a few mins.

The plan is to purchase a Visual fault locator which would enable us to quickly get a rough alignment of the fiber coupler. A local vendor is listed as a distributor for this product from JDSU.  Contact info:

DuVac Electronics (EDGE)
Tel: 626-796-3291
Email: jack@duvac.com
1759 E Colorado Blvd
Pasadena, CA 91106

 

ELOG V3.1.3-